OpenStudy (anonymous):

Daniela teaches Ballet. She charges a one-time processing fee of $3.00 and the cost of each class is shown below. Let x represent the number of classes and y represent the cost of classes. Based on this information, what will it cost, in dollars, to take 6 classes? (Do not put a $ in your answer.) Cost of classes (not including processing fee) Number of Classes, x 1 2 3 4 Cost of Classes, y $16.00 $32.00 $48.00 $64.00

OpenStudy (anonymous):

I know the answer is 89 but I dont know how to get it.

OpenStudy (anonymous):

Okay, find the relationship between x and y. 1 class is 16 dollars, right?

OpenStudy (anonymous):

so what's 6 x 16? 96.

OpenStudy (anonymous):

Oh ok thank you that makes alot of sense!

OpenStudy (anonymous):

Anytime :D

OpenStudy (anonymous):

wait but how do you get 89?

OpenStudy (anonymous):

It shouldn't be 89...it would actually be 99, since the three dollar processing fee.

OpenStudy (anonymous):

but the answer is 89... I have the answer sheet and everything.... (My assignment is to fully answer all of my wrong questions and my paper says 89 is the right answer)

OpenStudy (anonymous):

Let me look again.

OpenStudy (anonymous):

Could it be a typo maybe? Because 6*16=96, then plus 3 for the fee, makes 99.

OpenStudy (anonymous):

I really dont know my teacher probably isnt in right know (home schooled) so if I tried contacting her she probably wouldnt answer...

OpenStudy (anonymous):

Well, I'm 100% sure it's 99. Since that is the process to find the answer. I'd send her a message anyway, just to get reassurance.

OpenStudy (anonymous):

ok thank you could you help me with one more question? I forgot how to do the type of question

OpenStudy (anonymous):

Okay, sure.

OpenStudy (anonymous):

Okay What is the value for x in the proportion x/x+10 = 7/12 There are a bunch of these and I just need to know how to do them again.

OpenStudy (anonymous):

Okay, remember cross multiply?